OpenStudy (elianymendoza):

The first ionization energy, E, of a nitrogen atom is 2.32 aJ. What is the wavelength of light, in nm, that is just sufficient to ionize a nitrogen atom?

OpenStudy (aaronq):

Use the formula: \(E_{photon}=\dfrac{hc}{\lambda}\) where c is the speed of light, h plancks constant and \(\lambda\) the wavelength. Remember to use SI units.

OpenStudy (elianymendoza):

Ok, so I get the fact that: \[Ephoton=\frac{ (6.626 x 10^{-34}J*s)(2.9979*10^{8}m*s^{-1}) }{? }\] What I still don't understand is what wavelength is & where does the first ionization of nitrogen plays a role in this problem?

OpenStudy (aaronq):

The energy the photon needs to be is given in the question, 2.32 aJ. \(\sf 1~attojoule=1.0*10^{-18}J\) so, \(\sf E_{photon}=2.32~aJ=2.32*10^{-18}J\) wavelength is what you are solving for, they're asking for you to express it in nm. \(1 ~nm=1.0*10^{-9}~m\)